# Autonomous Navigation at the Nano-Scale: Algorithms, Architectures, and Constraints

### v1

@@ 1 Introduction
The remainder of this paper is organized as follows: Section 2 reviews the hardware platforms and low-power sensing architectures fundamental to the nano-scale. Section 3 dissects the “Onboard Autonomy Stack”, covering efficient machine learning perception, optimization-based planning, robust flight control, and decentralized swarm coordination, alongside the software toolchains required for deployment. Section 4 critically analyses the coupled physical, computational, and bandwidth constraints that define the system’s operational envelope. Section 5 evaluates the feasibility of these systems across diverse application domains, from industrial inspection to precision agriculture. Finally, Section 6 identifies persistent research gaps, and Section 7 concludes with a roadmap for future developments, advocating for a paradigm shift toward bio-inspired and neuromorphic architectures.
@@ 2.1 Processing Architectures
Table 1: Comparative Analysis of Compute Substrates: Standard vs. Nano-Scale
Device CPU / Arch Accelerator Memory Power Perf. Role
@@
STM32F4 [88] Cortex-M4 @ 168 MHz - 192 kB 100 mW <<0.5 GFLOPs Flight control
Teensy 4.0 [90] Cortex-M7 @ 600 MHz - 1 MB $\sim$100 mW - SNN research
GAP8 [29] RISC-V @ 250 MHz 8-core cluster 512 kB 86 mW 10 GMACs Edge AI
GAP9 [57] RISC-V @ 450 MHz 9-core cluster 1.5 MB 50 mW 3 GFLOPs AV / AI
Kraken [26] RISC-V @ 350 MHz SNN/ANN 1 MB 300 mW 3 GFLOPs Neuromorphic
Navion [92] Custom ASIC VIO engine On-chip 24 mW 171 FPS Visual odometry
Shaheen [97] RV64 @ 600 MHz 8-core cluster HyperRAM 200 mW 7.9 GFLOPs Linux + RTOS
Note: Top row (Jetson) provided as a baseline reference. Power figures represent typical operational envelopes. The Teensy 4.0 is included as a research platform for SNNs [90].
@@ 3.2.3 Optimization-Based Local Planning and Safety Filters
Research has prioritized adapting MPC specifically for constrained hardware by reducing solver complexity. Sampling-based methods, such as Model Predictive Path Integral (MPPI) control, offer robustness against non-smooth cost functions by sampling thousands of potential trajectories. However, standard MPPI implementations typically mandate GPU acceleration (e.g., NVIDIA Jetson) to achieve real-time performance, rendering them unsuitable for microcontroller-based nano-drones [pravitra2020ℒ]. To mitigate this, deep model predictive optimization has been proposed to learn the inner-loop optimizer of MPPI, achieving superior performance with fewer samples and reduced memory requirements [81]. Furthermore, to bridge the “Sim-to-Real” gap, L1-Adaptive control has been integrated with MPPI to compensate for both matched and unmatched uncertainties (such as aerodynamic drag) thereby recovering robustness without the heavy computational cost of full NMPC [pravitra2020ℒ]. Perhaps the most significant advancement for nano-UAVs is Tiny LB MPC [2], which explicitly targets the constraints of microcontrollers like the Teensy 4.0 (compatible with the Crazyflie 2.1). By exploiting the differential flatness of multirotor dynamics, this framework transforms the nonlinear optimal control problem into a convex second-order cone program. Crucially, it employs a custom Alternating Direction Method of Multipliers (ADMM) solver where expensive primal updates are pre-computed at a lower rate (10 Hz) using a linearized Gaussian Process. This architectural splitting allows the controller to execute at 100 Hz, outperforming standard flatness-based MPC by 23% in tracking error while remaining computationally feasible for edge deployment [2].
@@ 4 Key Challenges and Constraints
The transition of autonomous navigation from standard-sized quadrotors to nano-UAVs is not a mere scale reduction; it constitutes a fundamental redefinition of the operational envelope governed by cubic scaling laws. These scaling relations give rise to tight, interdependent trade-offs among flight endurance, sensing capability, and onboard computational resources. In this section we quantify those constraints and show how they motivate the specialised hardware architectures described in Section 4.1 and the aggressive algorithmic compression techniques presented in Section 3.
@@ 4.1 The Physics of Scaling: The Energy-Autonomy Trade-off
The primary constraint defining nano-UAV autonomy is the rigorous application of the square-cube law. As the characteristic length of the drone decreases to the sub-10 cm range [97], its volume and mass scale cubically, while the lifting surface area scales squarely. Consequently, the battery capacity drops precipitously; for a typical 27–33 g platform like the Crazyflie 2.x, the total power budget is approximately 7–10 W [72, 99]. However, aerodynamic lift consumes the vast majority of this budget; motors typically draw 95–96% of the total energy to maintain flight [72]. This physical reality leaves a minuscule power envelope for the avionics and onboard intelligence, typically bounded between 5% and 15% of the total system power [97, 63]. Quantitatively, this results in a strict compute budget of sub-100 mW to maximum 200–300 mW for all sensing, processing, and control tasks[97, 63]. This constraint creates a direct causal chain affecting autonomy: Every milliwatt consumed by the navigation engine subtracts directly from the flight time. For example, adding a dedicated visual navigation engine like the PULP-Shield (running the DroNet CNN) reduces the flight time by approximately 22% (from $\sim$440 s to $\sim$340 s) due to the combined effect of added payload weight (5 g) and electrical consumption [71, 21]. Consequently, the high-performance GPUs used in micro-aerial vehicles (e.g., NVIDIA Jetson, consuming >>10W) are physically impossible to deploy on nano-drones, necessitating the ultra-low-power microcontroller-based strategies discussed in Section .
@@ 4.2 The Deep Learning Memory Wall
This creates a dangerous “latency gap”: the drone must stabilize itself 30 to 80 times for every single visual update it receives. The causal chain of instability (Low Inertia $\to$ High-Frequency Dynamics (>>500Hz) $\to$ Perceptual Latency (>>50ms)) necessitates the hierarchical control architectures. High-level visual navigation engines cannot directly drive the motors; instead, they must generate setpoints (e.g., velocity or attitude commands) for a lower-level, high-frequency PID controller running on the MCU [72, 46]. Consequently, any interruption in the perception loop or communication link exceeding 20–40 ms can lead to catastrophic instability [98, 11]. However, emerging neuromorphic architectures are beginning to close this gap. Stroobants et al. [90] demonstrated that by replacing standard CNNs with SNNs for attitude control, the perception-actuation loop can operate at 500 Hz on a standard microcontroller, matching the frequency of the inner control loop and effectively eliminating the latency penalty associated with deep learning inference.
@@ 6.1 Fundamental Research Gaps
Despite the progress in specific application domains, the deployment of nano-UAVs is generally hindered by a series of interdependent physical and computational constraints that create a “Physics Gap” between required control frequency and available power.
@@ 6.2 Future Architectures and Recommendations
To bridge the fundamental performance gaps inherent to nano-UAVs, future research must pivot from component optimization to holistic architectural shifts in both silicon design and control paradigms. The current modular approach of stacking PCBs incurs unacceptable weight penalties, necessitating a convergence toward System-in-Package (SiP) solutions that integrate the flight controller, AI accelerator, and radio onto a single die to minimize interconnect weight and maximize performance-per-watt [97, 60]. Simultaneously, to close the latency gap, control loops should transition from discrete-time processing to event-based architectures; SNNs running on neuromorphic hardware can process sensor changes asynchronously with microsecond latency, matching the temporal dynamics of the platform [91]. Initial validation of this paradigm is already visible, with recent work successfully deploying end-to-end spiking control policies on commercial microcontrollers (Teensy 4.0) via imitation learning, proving that the benefits of event-based processing—specifically the reduction of floating-point overhead—can be realized even before the widespread availability of specialized neuromorphic hardware [90]. Finally, to address the “Sim-to-Real” gap, agents must possess the capacity for adaptation post-deployment; emerging frameworks like BackpropTools demonstrate the feasibility of running backpropagation directly on microcontrollers, allowing drones to fine-tune control policies in the field to compensate for battery degradation or motor damage [27].

### v2

@@ 1 Introduction
Distinct from prior literature, this review addresses the specific intersection of extreme SWaP constraints and onboard algorithmic autonomy. While Cai et al. [1] provided a foundational survey of small-scale platforms, their work primarily catalogs aerodynamic configurations and classical control strategies relevant to the pre-edge-AI era. Similarly, while Neuman et al. [18] effectively highlight the challenges of “Tiny Robot Learning,” their scope encompasses general micro-robotics (including terrestrial quadrupeds) with a primary emphasis on machine learning inference paradigms. In contrast, this review specifically targets the Nano-UAV domain, bridging the gap between hardware limitations and the full autonomy stack. We exclusively focus on the computational and algorithmic architectures required to close the perception-actuation loop entirely onboard under sub-100 mW constraints, moving beyond isolated subsystems to holistic autonomy.
1.1 Review Methodology and Scope
To provide a comprehensive analysis of the state-of-the-art, we surveyed literature from IEEE Xplore, ScienceDirect, and arXiv repositories, covering the period from 2014–2025. Special emphasis was placed on the post-2019 era, which marks the emergence of heterogeneous edge-AI accelerators (e.g., RISC-V GAP8) in the nano-robotics domain. The selection process prioritized experimental validity under strict Size, Weight, and Power (SWaP) constraints. Publications were selected based on three rigorous inclusion criteria:
  • Form Factor: We restricted our scope to “nano” class aerial platforms, defined here as vehicles with a total weight of <50<50 g and a diameter of <10<10 cm (e.g., Bitcraze Crazyflie, perceptually-enabled insect-scale robots). General Micro Aerial Vehicle (MAV) literature (>100>100 g) was excluded unless it introduced algorithmic primitives explicitly adapted for extreme quantization or low-inertia dynamics.
  • Onboard/Off-Board Autonomy: Priority was given to methodologies demonstrating fully onboard processing, where perception, state estimation, and control run locally on the microcontroller or/and edge accelerator. Approaches relying on external motion capture systems (e.g., Vicon/Optitrack) or offboard computation via radio streaming were excluded, except to serve as ground-truth baselines, due to their non-viability in real-world deployment.
  • Algorithmic Efficiency: The review targets works contributing to “TinyML,” spiking neural networks (SNNs), and lightweight control strategies (e.g., quantized DNNs, depth-based obstacle avoidance). We focused on algorithms capable of closing the perception-actuation loop within the milliwatt-range power envelopes of embedded SoCs, excluding heavy visual-inertial odometry (VIO) pipelines requiring desktop-class GPUs.
1.2 Organization of the Review
The remainder of this paper is organized as follows: Section  2 reviews the hardware platforms and low-power sensing architectures fundamental to the nano-scale. Section  3 dissects the “Onboard Autonomy Stack”, covering efficient machine learning perception, optimization-based planning, robust flight control, and decentralized swarm coordination, alongside the software toolchains required for deployment. Section  4 critically analyses the coupled physical, computational, and bandwidth constraints that define the system’s operational envelope. Section  5 evaluates the feasibility of these systems across diverse application domains, from industrial inspection to precision agriculture. Finally, Section  6 identifies persistent research gaps, and Section 7 concludes with a roadmap for future developments, advocating for a paradigm shift toward bio-inspired and neuromorphic architectures.
@@ 2.1 Processing Architectures
Table 1: Comparative Analysis of Compute Substrates: Standard vs. Nano-Scale (Normalized)
Device Architecture Power Peak Throughput Efficiency Memory Primary Role
(mW) (GOPS) (GOPS/W)
@@
STM32F4 [31] Cortex-M4 @ 168 MHz 100 <<0.5 <<5 192 kB Flight control
Teensy 4.0 [32] Cortex-M7 @ 600 MHz $\sim$100 1.2 $\sim$12 1 MB SNN research
GAP8 [33] RISC-V Cluster (8-core) 86 20.0b 232 512 kB Edge AI
GAP9 [34] RISC-V Cluster (9-core) 50 3.0c 60 1.5 MB AV / AI
Kraken [35] RISC-V + SNN Accel. 300 3.0 10 1 MB Neuromorphic
Navion [29] Custom ASIC (VIO) 24 Task-Specificd N/A On-chip Visual odometry
Shaheen [15] RV64 + RV32 Cluster 200 7.9c 39.5 HyperRAM Linux + RTOS
Note: Power figures represent typical operational envelopes. Efficiency is calculated as Peak Throughput / Power.
a Converted from 1.3 TFLOPs for unit consistency.
b Based on 10 GMACs ($1\text{ MAC}\approx 2\text{ OPS}$).
c Reported as GFLOPs (Floating Point).
d Throughput for ASICs is application-dependent (e.g., 171 FPS for Stereo-VIO [29]).
@@ 3.1.2 Optimization for the Edge: Quantization, Pruning, and Auto-Tiling
A critical distinction in this domain is the choice between Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ offers rapid deployment by calibrating pre-trained weights using a representative dataset, whereas QAT simulates quantization noise (e.g., rounding errors) during the training backward pass, forcing the network to adapt to lower precision at the cost of complex retraining. The evolution of the PULP-Dronet architecture illustrates the shifting preference between these paradigms as toolchains mature. The original implementation (V1) necessitated QAT to stabilize its 16-bit fixed-point control policy on the GAP8 SoC without accuracy loss [3, 67]. However, subsequent automated iterations (V2) transitioned to aggressive 8-bit PTQ leveraging the NEMO/DORY toolchain [66, 69]. This shift demonstrates that modern deployment flows can now preserve feature fidelity through advanced calibration statistics alone, eliminating the prohibitive computational overhead of retraining required by earlier architectures.
@@ 3.2 Navigation and Trajectory Generation
Navigation architectures for nano-UAVs generally fall into two distinct paradigms: Map-based (Modular) and Map-less (End-to-End). Map-based systems, such as NanoSLAM [55], decouple the autonomy stack into explicit mapping, planning, and control modules. This offers interpretability and safety guarantees through metric representations (e.g., occupancy grids) but incurs significant memory overhead to maintain state [55]. Conversely, Map-less approaches, exemplified by PULP-Dronet [3], utilize end-to-end learning to map raw sensor inputs directly to control actuation via neural networks. While this minimizes latency and memory footprint (ideal for SWaP-constrained nano-UAVs) it lacks the long-horizon planning capabilities and interpretability of modular stacks [3]. The following sections analyse how nano-UAVs navigate these trade-offs.
@@ 3.2.3 Optimization-Based Local Planning and Safety Filters
Research has prioritized adapting MPC specifically for constrained hardware by reducing solver complexity. Sampling-based methods, such as Model Predictive Path Integral (MPPI) control, offer robustness against non-smooth cost functions by sampling thousands of potential trajectories. However, standard MPPI implementations typically mandate GPU acceleration (e.g., NVIDIA Jetson) to achieve real-time performance, rendering them unsuitable for microcontroller-based nano-drones [83]. To mitigate this, deep model predictive optimization has been proposed to learn the inner-loop optimizer of MPPI, achieving superior performance with fewer samples and reduced memory requirements [84]. Furthermore, to bridge the “Sim-to-Real” gap, L1-Adaptive control has been integrated with MPPI to compensate for both matched and unmatched uncertainties (such as aerodynamic drag) thereby recovering robustness without the heavy computational cost of full NMPC [83]. Perhaps the most significant advancement for nano-UAVs is Tiny LB MPC [61], which explicitly targets the constraints of microcontrollers like the Teensy 4.0 (compatible with the Crazyflie 2.1). By exploiting the differential flatness of multirotor dynamics, this framework transforms the nonlinear optimal control problem into a convex second-order cone program. Crucially, it employs a custom Alternating Direction Method of Multipliers (ADMM) solver where expensive primal updates are pre-computed at a lower rate (10 Hz) using a linearized Gaussian Process. This architectural splitting allows the controller to execute at 100 Hz, outperforming standard flatness-based MPC by 23% in tracking error while remaining computationally feasible for edge deployment [61].
@@ 4 Key Challenges and Constraints
The transition of autonomous navigation from standard-sized quadrotors to nano-UAVs is not a mere scale reduction; it constitutes a fundamental redefinition of the operational envelope governed by cubic scaling laws. These scaling relations give rise to tight, interdependent trade-offs among flight endurance, sensing capability, and onboard computational resources. In this section we quantify those constraints and show how they motivate the specialised hardware architectures described in Section 2 and the aggressive algorithmic compression techniques presented in Section 3. We encapsulate these systemic bottlenecks in Figure 5, which illustrates the “scaling paradox”: a domain where geometric advantages are negated by viscous aerodynamics and energy density limits, forcing a hard compromise between flight time and intelligence.
@@ 4.1 The Physics of Scaling: The Energy-Autonomy Trade-off
While the Introduction defined the “Physics Gap” governing nano-scale flight, its practical consequence is a severe energy-autonomy trade-off. As the characteristic length of the drone decreases to the sub-10 cm range [15], the cubic scaling of mass versus the quadratic scaling of lifting surface area results in a precipitous drop in battery capacity (see Figure 5A). For a typical 27–33 g platform like the Crazyflie 2.x, the total power budget is approximately 7–10 W [25, 14]. However, aerodynamic lift consumes the vast majority of this budget; motors typically draw 95–96% of the total energy to maintain flight [25]. This physical reality leaves a minuscule power envelope for the avionics and onboard intelligence, typically bounded between 5% and 15% of the total system power [15, 69]. Quantitatively, this results in a strict compute budget of sub-100 mW to maximum 200–300 mW for all sensing, processing, and control tasks[15, 69]. This constraint creates a direct causal chain affecting autonomy: Every milliwatt consumed by the navigation engine subtracts directly from the flight time. For example, adding a dedicated visual navigation engine like the PULP-Shield (running the DroNet CNN) reduces the flight time by approximately 22% (from $\sim$440 s to $\sim$340 s) due to the combined effect of added payload weight (5 g) and electrical consumption [3, 45]. Consequently, the high-performance GPUs used in micro-aerial vehicles (e.g., NVIDIA Jetson, consuming >>10W) are physically impossible to deploy on nano-drones, necessitating the ultra-low-power microcontroller-based strategies discussed in Section  2.
@@ 4.2 The Deep Learning Memory Wall
This creates a dangerous “latency gap” (see Figure 5C): the drone must stabilize itself 30 to 80 times for every single visual update it receives. The causal chain of instability (Low Inertia $\to$ High-Frequency Dynamics (>>500Hz) $\to$ Perceptual Latency (>>50ms)) necessitates the hierarchical control architectures. High-level visual navigation engines cannot directly drive the motors; instead, they must generate setpoints (e.g., velocity or attitude commands) for a lower-level, high-frequency PID controller running on the MCU [25, 74]. Consequently, any interruption in the perception loop or communication link exceeding 20–40 ms can lead to catastrophic instability [57, 93]. However, emerging neuromorphic architectures are beginning to close this gap. Stroobants et al. [32] demonstrated that by replacing standard CNNs with SNNs for attitude control, the perception-actuation loop can operate at 500 Hz on a standard microcontroller, matching the frequency of the inner control loop and effectively eliminating the latency penalty associated with deep learning inference.
@@ 6.1 Fundamental Research Gaps
Despite the progress in specific application domains, the deployment of nano-UAVs is generally hindered by a series of interdependent physical and computational constraints that create a “Physics Gap” between required control frequency and available power. Table 2 provides a summary of these challenges and the corresponding architectural solutions.
@@
Table 2: Summary of Open Problems in Nano-UAV Autonomy. This table maps the fundamental constraints (Section 6) to required architectural shifts.
Challenge Current State (SoA) Remaining Gap Potential Solution
SWaP Modular PCBs; Avionics power <100<100 mW. The Physics Gap:
Energy budget precludes processors fast enough (>500>500 Hz) to handle low-inertia dynamics.
System-in-Package:
Single-die integration of flight controller, AI, and radio.
Latency CNNs on accelerators (e.g., GAP8) at 6–18 FPS. The Control Gap:
Drone requires 30–80 mechanical corrections per visual update; interruption leads to divergence.
Neuromorphic Eng.:
[…]
Memory Aggressive 8-bit quantization; Depth-wise convs. The Efficiency Gap:
Energy cost of data movement (RAM to CPU) exceeds computation cost.
Bio-inspired Arch.:
Computing-in-memory and sparse processing to bypass von Neumann limits.
Sim-to-Real Offline training; Static policies. The Adaptation Gap:
Inability to compensate for battery degradation or motor damage mid-flight.
On-Chip Learning:
TinyML frameworks (e.g., BackpropTools) for online fine-tuning.
Connectivity Off-loading compute to ground station. The Reliability Gap:
Bandwidth saturation and non-deterministic latency (>100>100 ms) violate stability criteria.
Indigenous Autonomy:
Full edge-processing without external dependence.
@@ 6.2 Future Architectures and Recommendations
To bridge the fundamental performance gaps inherent to nano-UAVs, future research must pivot from component optimization to holistic architectural shifts in both silicon design and control paradigms. The current modular approach of stacking PCBs incurs unacceptable weight penalties, necessitating a convergence toward System-in-Package (SiP) solutions that integrate the flight controller, AI accelerator, and radio onto a single die to minimize interconnect weight and maximize performance-per-watt [15, 16]. Simultaneously, to close the latency gap, control loops should transition from discrete-time processing to event-based architectures; SNNs running on neuromorphic hardware can process sensor changes asynchronously with microsecond latency, matching the temporal dynamics of the platform [28]. Initial validation of this paradigm is already visible, with recent work successfully deploying end-to-end spiking control policies on commercial microcontrollers (Teensy 4.0) via imitation learning, proving that the benefits of event-based processing (specifically the reduction of floating-point overhead) can be realized even before the widespread availability of specialized neuromorphic hardware [32]. Finally, to address the “Sim-to-Real” gap, agents must possess the capacity for adaptation post-deployment; emerging frameworks like BackpropTools demonstrate the feasibility of running backpropagation directly on microcontrollers, allowing drones to fine-tune control policies in the field to compensate for battery degradation or motor damage [95].